Overhead Variances, Four-Variance Analysis

Oerstman, Inc., uses a standard costing system and develops its overhead rates from the current annual budget. The budget is based on an expected annual output of 125,000 units requiring 500,000 direct labor hours. (Practical capacity is 520,000 hours.) Annual budgeted overhead costs total $840,000, of which $595,000 is fixed overhead. A total of 119,300 units using 498,000 direct labor hours were produced during the year. Actual variable overhead costs for the year were $262,000, and actual fixed overhead costs were $555,150.

Required:

1. Compute the fixed overhead spending and volume variances.

Fixed Overhead Spending Variance $ Favorable/Unfavorable
Fixed Overhead Volume Variance $ favorable/unfavorable

2. Compute the variable overhead spending and efficiency variances. Do not round intermediate calculations

Variable Overhead Spending Variance $ favorable./unfavorable
Variable Overhead Efficiency Variance $ favorable/unfavorable
Answer #1

Note: the following variances have been calculated based on budgeted hours, not practical capacity.

Requirement 1:

Fixed overhead spending variance

Fixed overhead spending variance is the difference between the budgeted fixed overhead and the actual fixed overhead. The variance is favourable when the budgeted fixed overhead is higher than the actual fixed overhead and is unfavourable when actual is higher than budgeted.

Fixed overhead spending variance= Budgeted fixed overhead –actual fixed overhead= 595000- 555150= $39850 Favourable

Fixed overhead volume variance

Fixed overhead volume variance is calculated by finding difference between the fixed overheads applied and the budgeted fixed overheads. When applied fixed overheads is higher than the budgeted fixed overheads, there is a favourable variance. If the opposite is the case, there is an unfavourable variance.

Fixed overhead volume variance= applied fixed overheads- budgeted fixed overheads= (Actual output- budgeted output) fixed overhead rate per unit

Fixed overhead rate per unit= budgeted fixed overhead/ budgeted units= 595000/125000= $4.76 per unit

Actual output= 119300 units

Budgeted output= 125000 units

Fixed overhead volume variance= (119300-125000) 4.76= $27132 unfavourable (because actual output is lower than budgeted, which results in a lower applied overhead compared to budgeted overhead)

Requirement 2:

Variable overhead spending variance

Variable overhead spending variance arises due the difference in standard variable overhead rate and the actual variable overhead rate. It is favourable when the actual rate is lower than the standard rate and unfavourable when the actual rate is higher than the standard rate.

Variable overhead spending variance= (standard rate- actual rate) actual hours= (Standard rate* actual hours)- (Actual rate * actual hours)= (Standard rate* actual hours)- actual variable overhead= (0.49*498000)- 262000= 244020- 262000= $17980 unfavourable

Standard rate= budgeted variable overhead/ budgeted hours= (840000-595000)/500000= $0.49 per hour

Variable overhead efficiency variance

This variance arises due the over efficiency or under efficiency in the application and use of variable overheads by the manufacturing unit (or factory). When the standard hours for the actual output are higher than the actual hours, there is a favourable variance. If the opposite is the case, it us unfavourable.

Variable OH efficiency variance= (Standard hours-Actual hours) standard rate=

(477200-498000) 0.49= $10192 unfavourable

Standard hours for actual output= (500000/ 125000)* 119300= 477200 hours